OpenStudy (anonymous):

tx''-(t+1)x'+x=0 t>0 f(t)=e^t need help finding my mistake! Please In Problems 45 through 48, a differential equation and a non-trivial solution f are given. Find a second linearly independent solution using reduction of order.

OpenStudy (anonymous):

http://prntscr.com/1i4ms4

OpenStudy (anonymous):

@primeralph think you could help me out on this one?

OpenStudy (anonymous):

i was able to do the last problem that was similar to this

OpenStudy (anonymous):

would you like to see the last problem i worked?

OpenStudy (primeralph):

Hold on.

OpenStudy (primeralph):

@KingGeorge

OpenStudy (primeralph):

@satellite73

OpenStudy (anonymous):

should i close the question? is anyone here?

OpenStudy (anonymous):

@dumbcow wanna take a look at this before i close it?

OpenStudy (psymon):

I'm insanely new to this kind of stuff, but I'm kinda just reading it over for the hell of it x_x

OpenStudy (anonymous):

lol

OpenStudy (psymon):

I'll have to do it in a month anyway myself, so might as well see what Im in for.

OpenStudy (anonymous):

i feel like a zombie all ive been doing all day everyday for the last 3 weeks is differential equations. been neglecting phys 2 which sucks but this class is a prereq for two im taking in the fall so i NEED TO pass it

OpenStudy (psymon):

Yeah, I hear ya. I'm taking it starting at the end of August. I'm reluctant to buy out textbook yet since apparently it's crap. But I have been looking over some of the things that will be in the class and I recognized the reduction of order so I thought Id go look.

OpenStudy (anonymous):

im gonna close the question. thanks for looking everyone!

OpenStudy (psymon):

Out of curiosity, did you actually get an answer when trying to solve?

OpenStudy (anonymous):

no i couldn't get past the part of substituting and getting a v'' and v' i got two v's for some reason. i did cancel the v though which was good but the two v' throw me off

OpenStudy (psymon):

I see. I got past that part actually. I may nothave a correct answer, but I'm on to something. Almost done and I can give ya an answer, which may be a million miles off, but an answer.

OpenStudy (anonymous):

ok

OpenStudy (psymon):

Hmm....Probably a mistake, but somehow I ended up with something where nearly everything cancelled out. No idea if that's an automatic sign of a wrong answer, but the second part of the general solution I got was simply C2(-t-1). After all the substitutions, things just canceled out. Perhaps the linearly independent part that I do not know about throws a monkey wrench into it. But I can maybe show where I got the v'' and v' to work. Although dumbcow may have all your answers, too, haha.

OpenStudy (psymon):

*Had fun trying no matter what* :)

OpenStudy (dumbcow):

sorry i was looking up some stuff which method did you use? here is a good site to use http://tutorial.math.lamar.edu/Classes/DE/LaplaceTransforms.aspx also in case you dont have solution: http://www.wolframalpha.com/input/?i=tx%28t%29%27%27-%28t%2B1%29x%28t%29%27%2Bx%28t%29%3D0 wolfram uses LaPlace transforms

OpenStudy (anonymous):

yeah i get excited if i even get close to the ans cause then i know its just a small mistake somewhere but unfortunately this is an even problem so i don't know the answer.

OpenStudy (psymon):

Thats actually where I looked. I tried following along with the reduction of order section. It at least led me to the point where I could even get an answer.

OpenStudy (anonymous):

i tried using reduction of order

OpenStudy (anonymous):

@rperez36 What is the class this work is from called?

OpenStudy (anonymous):

differential equations

OpenStudy (anonymous):

hello are you still having trouble?

OpenStudy (anonymous):

so where is my mistake? i checked over and over my product rule and distribution but couldn't see why i have two v'

OpenStudy (anonymous):

yes still having trouble lol

OpenStudy (psymon):

I had two v' as well if it makes ya feel better, lol.

OpenStudy (anonymous):

@rperez36 In the 4th line of the picture, where did you get it from?

OpenStudy (anonymous):

i think we are only supposed to have a v'' and v'

OpenStudy (ybarrap):

i applied red of order using your equations and found that \(v=\int\limits_{}^{}\frac{ e ^{\frac{ -1 }{ t }} }{ t }dt\)

OpenStudy (psymon):

Thats what id be unsure of. I just did reduction of order anyway, even though I had two v'. I don't know the type of problems well enough to know if that's even legal xD

OpenStudy (anonymous):

i did v times the known solution e^x which is what i did for the last problem and i was able to work it

OpenStudy (ybarrap):

I created a variable \(w=v'\) and solved the 1st order, separable equation that resulted

OpenStudy (psymon):

Yeah, that was the step I basically did after I had the two v's. I just did the w' = v'' and w=v'

OpenStudy (anonymous):

can we have two v' ?? if so then i shoulve just continued

OpenStudy (psymon):

I continued just for the hell of it to see if I'd get anything that'd look logical.

OpenStudy (anonymous):

tx''-(t+1)x'+x=0 t>0 f(t)=e^t need help finding my mistake! Please In Problems 45 through 48, a differential equation and a non-trivial solution f are given. Find a second linearly independent solution using reduction of order. For reduction of order, assume there exists another solution \(x(t)=u(t)e^t\) thus \(x'(t)=u'(t)e^t+u(t)e^t,x''(t)=u''(t)e^t+2u'(t)e^t+u(t)e^t\). Substituting in we get:$$t(u''e^t+2u'e^t+ue^t)-(t+1)(u'e^t+ue^t)+ue^t=0\\tu''e^t+2tu'e^t+tue^t-tu'e^t-tue^t-u'e^t-ue^t+ue^t=0\\tu''e^t+tu'e^t-u'e^t=0\\tu''+(t-1)u'=0$$

OpenStudy (ybarrap):

after doing this, I got \(\int\limits_{}^{}\frac{ dw }{ w }=\int\limits_{}^{}\frac{ \frac{ 1 }{ t } -1}{ t }dt\)

OpenStudy (psymon):

I got something pretty damn close to what ybarrap got it seems.

OpenStudy (psymon):

If I knew how to understand when math is written that way then maybe I could see my mistake, too.

OpenStudy (ybarrap):

@oldrin.bataku completed it. Looks good

OpenStudy (anonymous):

what happened to the e^t from (t-1)u'

OpenStudy (anonymous):

@rperez36 i divided thru by \(e^t\) (this is safe because \(e^t\ne0\))

OpenStudy (anonymous):

oh nice! i didn't even realize they were all gone. Nice job!

OpenStudy (anonymous):

i had a feeling there was a "trick" involved lol

OpenStudy (psymon):

Since I'm new to this, I apologize for the stupid question, but once you get to v = te^(-t), you don't have to integrate it back into terms of u? Where you were before the reduction oforder I mean.

OpenStudy (psymon):

I made it that far and thought I had to integrate it back.

OpenStudy (anonymous):

well you substitute a variable for v' then make variable' equal v'' to make it a first order equation seperate then solve for variable then replace back with v' and integrate from there if that makes any sense. i need to get to bed have class in the am

OpenStudy (psymon):

Night then.

OpenStudy (anonymous):

night everyone!

OpenStudy (anonymous):

@Psymon you are totally correct I made a very dumb error. sorry @rperez36

OpenStudy (anonymous):

Let \(v=u'\) to reduce it to a first-order separable ODE:$$tv'+(t-1)v=0\\tv'=(1-t)v\\\frac1vdv=\frac{1-t}tdt\\\int\frac1vdv=\int\left(\frac1t-1\right)dt\\\log v=\log t-t\\v=e^{\log t}e^{-t}=te^{-t}$$Switching back to \(u\) we have:$$u=\int te^{-t}\,dt=-te^{-t}+\int e^{-t}\,dt=-te^{-t}-e^{-t}=-e^{-t}(t+1)$$... hence our solution is given by \(x=ue^t=-(t+1)\) hence our second solution is merely \(-(t+1)\) (or more simply \(t+1\) as the coefficient does not matter -- our equation is linear)

OpenStudy (psymon):

Hm...I somehow got it right. Ican sleep happy now. Thanks for finishing that problem, I wanted to see the answer as well ^_^